# CloudFix: Automated Policy Repair for Cloud Access Control Policies Using Large Language Models

### v1

@@ VI Discussion and Threat to Validity
In this study, our main contribution is to leverage LLMs for policy repair and showing that Fault Localization-guided prompt enables LLMs to repair with improved accuracy and time, particularly for moderately complex request sizes (RQ1, RQ2). Our results show that the choice of LLMs plays a key part in performance, as demonstrated in RQ3, with CodeLlama-7B-Instruct achieving the highest complete repair. Results obtained in RQ4 show that LLMs can be used to repair policies that generalize to unseen, semantically similar requests. We further showed that LLMs can be used to synthesize access requests, which can be used to repair faulty access control policies (RQ5). We noticed the baseline repair took significantly more time for all repairs than FL, as shown in Table IV. This was often because the Baseline was attempting to enumerate the requests in an attempt to repair the policy, which often resulted in generating a failed policy (<80%). Our study did not use a timeout because we planned to focus on fairly evaluating the repair capability of our framework. Finally, our study relied on access control policies we curated from AWS re:Post, and algorithmically generated requests based on the structure of these policies; however, these policies may not fully capture the nuanced access control policies in real-world. The generalizability of our findings is constrained by the size of requests. We observed a degradation of repair accuracy and an increment of the overall repair time as the number of requests went up to 50, where FL generated only 7 complete repairs. Future work should consider techniques to maintain better accuracy for higher request sizes.
Beyond repair accuracy of our framework, the practicality of CloudFix hinges on its computational efficiency. Our analysis in Table IV shows that the FL-guided approach is significantly time-efficient than the baseline counterpart. However, more than 99% of the repair time was LLM inference time; therefore, future work should investigate the trade-off between accuracy and efficiency in model selection.
@@ VII Related Work
Access Control Policy Repair: Significant research has been conducted towards verifying and maintaining correct access control policies [10.5555/3161326, DetecingResolvingPolicyMisconfig, 10.1145/1978942.1979243, KERN2022103301, Jayaraman2014AutomatedAA, Son2013FixMU, UsabilityChallengesAccessControl, UsabilityConsiderationsAccessControl, EffectsPolicyConflictResoution]. Numerous access control policy repair techniques based on formal methods have been proposed to fix policies to adhere to the principle of least permissiveness. Prior work by [xu2016towards] introduces a fault localization and mutation-based policy repair method to debug and XACML access control policies. D’Antoni et al. [d2024automatically] proposed a synthesis-based framework to automatically reduce permissiveness in (AWS) IAM policies by leveraging access logs. Similarly, [peng2017towards] introduced fault localization and mutation-based policy repair techniques to produce a list of suspicious elements in XACML policies by correlating the test results and the test coverage information. They used mutation-based policy repair searches for fixes by mutating suspicious policy elements with predefined mutation operators. Eiers et al. [eiers2023quantitative] proposed a quantitative symbolic analysis framework to automatically repair overly permissive cloud access control policies. They encoded policy semantics as SMT formulas and used model counting to quantify and iteratively reduce the permissiveness of access control policies. However, their approach cannot repair semantically faulty policies.
@@
LLM-Based Automated Program Repair: Recent methods have shown promising results that LLMs can be effectively applied to automated program repair [zhang2024systematic][li2025hybrid][hossain2024deep][xia2023automated][campos2025empirical]. These approaches frame bug fixing as a code generation problem. A recent work closely related to ours, [orvalho2025counterexample], introduces a formal method based on counterexamples to guide LLMs to provide more accurate program repairs. However, to the best of our knowledge, there is no prior work that has explored automated access control policy repair with LLMs guided by a formal method-based fault localization.
VIII Conclusion
@@
In this work, we introduced CloudFix, the first automated framework to integrate LLMs with Formal Methods for access control policy repair. Manually repairing policies is error-prone and time-consuming. We address this problem by developing a framework to automatically create policy repairs with LLMs and by enhancing their reasoning capabilities with targeted repairs, enabled by Fault Localization. Our evaluation on a curated dataset of 282 real-world AWS policies showed that LLMs enhanced with Fault Localization can be used to repair cloud access control policies with significantly improved accuracy than the baseline approach.
@@ IX Data Availability
The policy dataset supporting this paper’s findings can be found here: https://doi.org/10.5281/zenodo.17373778

### v2

@@ VI Discussion and Threat to Validity
We provide a qualitative analysis and discuss our findings. To understand the behavior of CloudFix at scale, we analyzed the failure and success modes of CodeLlama-7B-Instruct. The 7 successful repairs at request size 50 all involved single-statement policies amenable to straightforward additive repair, with 86% completing within 2 iterations. The model consistently avoided overly permissive wildcards, preferring to enumerate specific requests. For moderate (80–99%) repairs, the LLM generated syntactically valid policies but introduced semantic errors such as omitting explicit deny statements. For failed (<<80%) repairs, we observed more severe failures: condition omission (e.g., dropping IpAddress constraints), entity hallucination (substituting arn:aws:iam::123456789012:root for the actual principal), and syntactic degeneration where 46% of policies showed no improvement due to repeated invalid JSON generation until token exhaustion. These behaviors align with known LLM tendencies to revert to high-frequency training patterns on out-of-distribution inputs [12].
Building on these observations, our main contribution is to show that fault localization guided prompting enables LLMs to repair policies with improved accuracy and time, particularly for moderately complex request sizes (RQ1, RQ2). The choice of LLM significantly impacts performance (RQ3), with CodeLlama-7B-Instruct achieving the highest complete repair rate, likely due to its code-focused pretraining providing familiarity with structured JSON, though we cannot isolate specific contributing factors. Results from RQ4 show that repairs generalize to unseen, semantically similar requests, and RQ5 demonstrates that LLM-synthesized requests from natural language can drive effective repair.
The baseline took significantly more time across all request sizes (Table IV) as it enumerated requests individually, resulting in failure. We did not impose a timeout to ensure fair evaluation. Our dataset relies on policies from AWS re:Post with algorithmically generated requests, which may not fully capture real-world complexity. Additionally, generalizability is constrained by request size: accuracy degraded as request size grew to 50, with Fault Localization generating only 7 complete repairs.
Beyond accuracy, the practicality of CloudFix depends on computational efficiency. The Fault Localization guided approach is significantly more time-efficient than the baseline, though >>99% of repair time is spent on LLM inference. We restricted evaluation to smaller open-source models due to hardware constraints and to demonstrate that privacy-preserving policy repair is achievable without proprietary models. Our results show that combining post-hoc verification with Fault Localization provides formal correctness guarantees for small to medium request sizes while maintaining policy confidentiality.
@@ VII Related Work
Access Control Policy Repair: Significant research has been conducted towards verifying and maintaining correct access control policies [18, 1, 23, 15, 13, 25, 22, 2, 21]. Numerous access control policy repair techniques based on formal methods have been proposed to fix policies to adhere to the principle of least permissiveness. Prior work by [29] introduces a fault localization and mutation-based policy repair method to debug XACML access control policies. D’Antoni et al. [5] proposed a synthesis-based framework to automatically reduce permissiveness in (AWS) IAM policies by leveraging access logs. Similarly, [20] introduced fault localization and mutation-based policy repair techniques to produce a list of suspicious elements in XACML policies by correlating the test results and the test coverage information. They used mutation-based policy repair searches for fixes by mutating suspicious policy elements with predefined mutation operators. Eiers et al. [7] proposed a quantitative symbolic analysis framework to automatically repair overly permissive cloud access control policies. They encoded policy semantics as SMT formulas and used model counting to quantify and iteratively reduce the permissiveness of access control policies. However, their approach cannot repair semantically faulty policies. Additionally, traditional symbolic repair tools, such as those by [4] and [9], are limited to specific, hard-coded repair patterns and will not generalize to semantically complex policies. Conversely, our approach leverages the pattern recognition capabilities of LLMs to understand the semantic relationship among large and high-dimensional policies and requests to synthesize complex repairs beyond simple mutations.
@@
LLM-Based Automated Program Repair: Recent methods have shown promising results that LLMs can be effectively applied to automated program repair [30, 17, 10, 28, 3]. These approaches frame bug fixing as a code generation problem. A recent work closely related to ours, [19], introduces a formal method based on counterexamples to guide LLMs to provide more accurate program repairs. However, to the best of our knowledge, there is no prior work that has explored automated access control policy repair with LLMs guided by a fault localization.
VIII Conclusion and Future Work
@@
In this work, we introduced CloudFix, the first automated framework to integrate LLMs with Formal Methods for access control policy repair. Manually repairing policies is time-consuming and prone to errors. We address this problem by developing a framework to automatically create policy repairs with LLMs and by enhancing their reasoning capabilities with targeted repairs, enabled by Fault Localization guidance. Our evaluation of a curated dataset of 282 real-world AWS policies showed that LLMs guided by Fault Localization can be used to repair cloud access control policies with significantly improved accuracy compared to the baseline approach. Future work could explore larger open-source models with greater context capacity, and apply supervised fine-tuning or reinforcement learning with verifiable rewards—the low computational cost of the SMT verifier makes it a practical reward signal.
@@ IX Data Availability
The policy dataset and accompanying source code used in this study are publicly available at https://github.com/bethelhall/fixmypolicy.